OpenStudy (anonymous):

The strength of a rectangular beam varies jointly as its width and the square of its depth. If the strength of a beam 12 inches wide by 22 inches deep is 3400 pounds per square inch, what is the strength of a beam 9 inches wide and 18 inches deep?

OpenStudy (anonymous):

let strength of beam = s The strength varies jointly as its width and the square of its depth. s = kwd^2 where k =constant of proportionality w=width d=depth 1)s=3400 pounds w=12 inches d=22 inches k=?? plug values in formula s = kwd^2 3400=k*12*22*22 k=0.585 2) s=?? w=9 inches d=18 inches k=0.585 plug in formula s = kwd^2 s=0.585*9*18*18 s=1705.86~1706 pounds